OpenStudy (anonymous):

If a-6, a, a+5 is a geometric sequence, then the common ratio of the sequence is ....?

ganeshie8 (ganeshie8):

take a proportion and solve a

ganeshie8 (ganeshie8):

\(\large \frac{a}{a-6} = \frac{a+5}{a}\)

ganeshie8 (ganeshie8):

or another easy way wud be to assume r is the common ratio and setup equations :- (a-6)r = a -------------(1) (a-6)r^2 = a+5---------(2)

OpenStudy (anonymous):

a=-30

ganeshie8 (ganeshie8):

take (2)/(1) and give it a try, which ever way u find it easy..

ganeshie8 (ganeshie8):

yes a = -30

OpenStudy (anonymous):

r=5/6?

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

oic

OpenStudy (anonymous):

thx

ganeshie8 (ganeshie8):

gw !